OpenStudy (anonymous):

Express the exact value of csc 60 degrees, with a rational denominator.

OpenStudy (anonymous):

Can someone plz help me with this

OpenStudy (asnaseer):

csc(60)=1/sin(60) sin(60) has a well known value - do you know the value of sin(60)?

OpenStudy (anonymous):

yes it is radical 3/2

OpenStudy (asnaseer):

yes, so we get:\[\csc(60)=\frac{2}{\sqrt{3}}\]

OpenStudy (asnaseer):

now just multiply the numerator and denominator by \(\sqrt{3}\)

OpenStudy (asnaseer):

that will get rid of the radical from the denominator

OpenStudy (anonymous):

ok thank so much:), It makes perfect sense:)

OpenStudy (asnaseer):

yw